# Fast Direct: Query-Efficient Online Black-box Guidance for Diffusion-model Target Generation

### v4

@@ Appendix B Ablation Study
We perform an ablation study on the target image generation Task-1. For step size analysis, we fix the batch size as $32$, then perform experiment with different step size $\alpha=\{20,40,80,160,320\}$; for batch size analysis, we fix the step size as $80$, then perform experiment with different batch size $B=\{4,8,16,32,64\}$. Additionally, we report the run time for different batch size. We report the result in Fig. 8.
We can observe that the performance steadily increase for any step size and any batch size, suggests that our algorithm is not sensitive to the hyper-parameters settings. The run time scales linearly with the batch size. In our target image generation experiments in Section 4.1, as the batch size is set as $32$, each experiment takes approximately $6.4$ hours to process $32$ images in parallel.
@@ Appendix D Baselines Details
For DDPO, DPOK and D3PO, we fine-tune the model to maximize the Gemini rating. We fine-tune the model with $50$ epochs; each epoch utilizes one batch query for model updating. For DDPO and DPOK, we set the batch size to $32$; for D3PO, it requires a relative reward, so we doubled the batch size to $64$. [^7] For D3PO, to evaluate $64$ images only requires $32$ queries calls, see Appendix G. As these RL-based methods require closed-form expression of the logarithm probabilities, we follow their official implementation to use the DDIMScheduler (Song et al., 2020) sampler.
For DNO, we optimize the noise sequence w.r.t the Gemini rating. Recall that the Gemini is black-box, so we use the non-differentiable mode, with the number of samples for gradient approximation set as $32$, and optimize for $50$ iterations; each iteration utilizes one batch query for updating. However, each experiment of DNO only produces one image, where the score is highly dependent on the initial prior. Thus, for a more accurate evaluation, we run 16 independent experiments to generate 16 images and report the average results. Note that this requires $16\times 50=800$ batch queries for each task, which is $16$ times compared with our Fast Direct and other baselines.
@@ Appendix F Overview of Related Works
We present the summary of related works according to its category and supported problem scenarios in Table 2.
@@ Appendix G Experiment Details
We use the following query question to the Gemini:
@@ Appendix H Molecules Generation Experiment Results
We report the average Vina score in molecules generation task in Fig. 10.
Figure 10: The Vina score (lower is better) of the generated molecules for each number batch queries on the six protein receptors.
@@ Appendix I More Experiment Results
We report the accumulated objective values over number of batch queries for images task in Fig. 11, and for molecules task in Fig. 12.
Figure 11: The accumulated Gemini rating (from 1 to 5, higher is better) over number of batch queries. The accumulated plot displays the maximum Gemini rating achieved up to each number of batch queries.
Figure 12: The accumulated Vina score (lower is better) over number of batch queries. The accumulated plot displays the minimum Vina score achieved up to each number of batch queries.
@@
Table 3 shows the list the complete prompts used for each task in our image generation experiment. From Fig. 13 to Fig. 23, we present the generated images of each algorithms over each number of batch queries for prompt 2 to 12. From Fig. 24 to Fig. 34, we showcase the 32 randomly generated by pre-trained model followed by the resultant images guided by the Fast Direct for prompt 2 to 12.
@@
Figure 13: The generated images over each number of batch queries on the prompt ”traffic-light”, extra batch queries budgets (until 500) are given to the baselines methods for demonstration.
Figure 14: The generated images over each number of batch queries on the prompt ”apple”, extra batch queries budgets (until 500) are given to the baselines methods for demonstration.
Figure 15: The generated images over each number of batch queries on the prompt ”cyber-dog” for each algorithms.
Figure 16: The generated images over each number of batch queries on the prompt ”puppy-nose” for each algorithms.
Figure 17: The generated images over each number of batch queries on the prompt ”robot-plant” for each algorithms.
Figure 18: The generated images over each number of batch queries on the prompt ”ocean” for each algorithms.
Figure 19: The generated images over each number of batch queries on the prompt ”sand-glass” for each algorithms.
Figure 20: The generated images over each number of batch queries on the prompt ”penguin” for each algorithms.
Figure 21: The generated images over each number of batch queries on the prompt ”basket” for each algorithms.
Figure 22: The generated images over each number of batch queries on the prompt ”ice-cube” for each algorithms.
Figure 23: The generated images over each number of batch queries on the prompt ”cat-butterfly” for each algorithms.
@@
Figure 24: The 32 randomly generated image for the prompt ”traffic-light” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 25: The 32 randomly generated image for the prompt ”apple” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 26: The 32 randomly generated image for the prompt ”cyber-dog” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 27: The 32 randomly generated image for the prompt ”puppy-nose” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 28: The 32 randomly generated image for the prompt ”robot-plant” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 29: The 32 randomly generated image for the prompt ”ocean” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 30: The 32 randomly generated image for the prompt ”sand-glass” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 31: The 32 randomly generated image for the prompt ”penguin” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 32: The 32 randomly generated image for the prompt ”basket” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 33: The 32 randomly generated image for the prompt ”ice-cube” guided by Fast Direct by utilizing 50 batch queries budget.
@@
Figure 34: The 32 randomly generated image for the prompt ”cat-butterfly” guided by Fast Direct by utilizing 50 batch queries budget.

### v5

@@ Appendix B Ablation Study
We perform an ablation study on the target image generation Task-1. For step size analysis, we fix the batch size as $32$, then perform experiment with different step size $\alpha=\{20,40,80,160,320\}$; for batch size analysis, we fix the step size as $80$, then perform experiment with different batch size $B=\{4,8,16,32,64\}$. Additionally, we report the run time for different batch sizes. We report the result in Fig. 8.
We can observe that the performance steadily increases for any step size and any batch size, which suggests that our algorithm is not sensitive to the hyperparameter settings. The run time scales linearly with the batch size. In our target image generation experiments in Section 4.1, as the batch size is set as $32$, each experiment takes approximately $6.4$ hours to process $32$ images in parallel.
@@ Appendix D Baselines Details
For DDPO, DPOK, and D3PO, we fine-tune the model to maximize the Gemini rating. We fine-tune the model with $50$ epochs; each epoch utilizes one batch query for model updating. For DDPO and DPOK, we set the batch size to $32$; for D3PO, it requires a relative reward, so we doubled the batch size to $64$. [^7] For D3PO, to evaluate $64$ images only requires $32$ queries calls, see Appendix G. As these RL-based methods require closed-form expression of the logarithm probabilities, we follow their official implementation to use the DDIMScheduler (Song et al., 2020) sampler.
For DNO, we optimize the noise sequence w.r.t the Gemini rating. Recall that the Gemini is a black-box, so we use the non-differentiable mode, with the number of samples for gradient approximation set as $32$, and optimize for $50$ iterations; each iteration utilizes one batch query for updating. However, each experiment of DNO only produces one image, where the score is highly dependent on the initial prior. Thus, for a more accurate evaluation, we run 16 independent experiments to generate 16 images and report the average results. Note that this requires $16\times 50=800$ batch queries for each task, which is $16$ times compared with our Fast Direct and other baselines.
@@ Appendix F Overview of Related Works
We present the summary of related works according to their category and supported problem scenarios in Table 2.
@@ Appendix G Experiment Details
We use the following query question for the Gemini:
@@ Appendix H Molecules Generation Experiment Results
We report the average Vina score in the molecules generation task in Fig. 10.
Figure 10: The Vina score (lower is better) of the generated molecules for each number of batch queries on the six protein receptors.
@@ Appendix I More Experiment Results
We report the accumulated objective values over the number of batch queries for the images task in Fig. 11, and for the molecules task in Fig. 12.
Figure 11: The accumulated Gemini rating (from 1 to 5, higher is better) over the number of batch queries. The accumulated plot displays the maximum Gemini rating achieved for each number of batch queries.
Figure 12: The accumulated Vina score (lower is better) over the number of batch queries. The accumulated plot displays the minimum Vina score achieved up to each number of batch queries.
@@
Table 3 shows the list the complete prompts used for each task in our image generation experiment. From Fig. 13 to Fig. 23, we present the generated images of each algorithm over each number of batch queries for prompts 2 to 12. From Fig. 24 to Fig. 34, we showcase the 32 randomly generated by the pre-trained model, followed by the resultant images guided by the Fast Direct for prompts 2 to 12.
@@
Figure 13: The generated images over each number of batch queries on the prompt ”traffic-light”, extra batch query budgets (until 500) are given to the baseline methods for demonstration.
Figure 14: The generated images over each number of batch queries on the prompt ”apple”, extra batch query budgets (until 500) are given to the baseline methods for demonstration.
Figure 15: The generated images over each number of batch queries on the prompt ”cyber-dog” for each algorithm.
Figure 16: The generated images over each number of batch queries on the prompt ”puppy-nose” for each algorithm.
Figure 17: The generated images over each number of batch queries on the prompt ”robot-plant” for each algorithm.
Figure 18: The generated images over each number of batch queries on the prompt ”ocean” for each algorithm.
Figure 19: The generated images over each number of batch queries on the prompt ”sand-glass” for each algorithm.
Figure 20: The generated images over each number of batch queries on the prompt ”penguin” for each algorithm.
Figure 21: The generated images over each number of batch queries on the prompt ”basket” for each algorithm.
Figure 22: The generated images over each number of batch queries on the prompt ”ice-cube” for each algorithm.
Figure 23: The generated images over each number of batch queries on the prompt ”cat-butterfly” for each algorithm.
@@
Figure 24: The 32 randomly generated images for the prompt ”traffic-light” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 25: The 32 randomly generated images for the prompt ”apple” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 26: The 32 randomly generated images for the prompt ”cyber-dog” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 27: The 32 randomly generated images for the prompt ”puppy-nose” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 28: The 32 randomly generated images for the prompt ”robot-plant” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 29: The 32 randomly generated images for the prompt ”ocean” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 30: The 32 randomly generated images for the prompt ”sand-glass” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 31: The 32 randomly generated images for the prompt ”penguin” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 32: The 32 randomly generated images for the prompt ”basket” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 33: The 32 randomly generated images for the prompt ”ice-cube” guided by Fast Direct by utilizing 50 batch query budget.
@@
Figure 34: The 32 randomly generated images for the prompt ”cat-butterfly” guided by Fast Direct by utilizing 50 batch query budget.